OpenStudy (anonymous):

How do I rewrite sin³x so that it does not contain powers greater than 1?

OpenStudy (lgbasallote):

uhmm factor it out?

Parth (parthkohli):

\[\sin (x) \sin (x) \sin (x)\]Is that what you mean?

OpenStudy (anonymous):

I'd do it like this \[\sin (x)\sin ^{2}(x)\]

zepdrix (zepdrix):

 Is this what you were trying to do? :o